A teacher's income from his investment in balloon bank and monthly wages is $100000. he doubles his investment and also gets an increase of 50% in wages and his income increases to $160000. what was his original income separately in terms of dividends and wages
Found 3 solutions by mananth, ikleyn, MathTherapy:
Answer by mananth(16946)


A teacher's income from his investment in balloon bank and monthly wages is $100000. he doubles his investment and also gets an increase of 50% in wages and his income increases to $160000. what was his original income separately in terms of dividends and wages
investment in balloon bank be x
Income in wages be y
Total income 100000
x+y=100000-------------(1)

doubles his investment =2x
increase in wages = 1.5y
2x+1.5y 160000-------------------(2)
multiply (1) by (2) qnd subtract (2)
2x+2y=200000
2x+1.5y = 160,000
0.5y = 40000
y =80000
x=20000

Answer by ikleyn(52781)

.


                    To the dear visitor/creator of this problem :



I do not think that it is correct here, in this problem, to talk about "monthly wages".


Change  "monthly wages"  to  "annual wages"  in the problem's description,  to keep the problem in more realistic/reasonable frame.



Answer by MathTherapy(10552)

A teacher's income from his investment in balloon bank and monthly wages is $100000. he doubles his investment and also gets an increase of 50% in wages and his income increases to $160000. what was his original income separately in terms of dividends and wages
Let original income from investment be I
Then income from wages = 100,000 - I
After doubling his investment income, it becomes 2I and after a 50% increase in wages, it becomes 1.5(100,000 - I)
We then get: 2I + 1.5(100,000 - I) = 160,000
2I + 150,000 - 1.5I = 160,000
2I - 1.5I = 160,000 - 150,000
.5I = 10,000
Original income from investments, or
Now, you should be able to find the income from wages.